OpenStudy (anonymous):

NEED HELP FAST WILL FAN AND MEDAL! Is triangle ABC a right triangle? Why or why not?

OpenStudy (anonymous):

No because a right triangle has 90 degrees in it but since none of them have the measurement ABC is not a right triangle

OpenStudy (anonymous):

how do you post pictures

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

And what would AC have to be to make it a right triangle?

OpenStudy (anonymous):

1 thing ...

OpenStudy (anonymous):

To solve an SSS triangle: use The Law of Cosines first to calculate one of the angles. then use The Law of Cosines again to find another angle. and finally use angles of a triangle add to 180° to find the last angle.

OpenStudy (anonymous):

soh cah toa ...

OpenStudy (anonymous):

agree to @jakashaka123

OpenStudy (anonymous):

i remember it as some old hippie caught another hippie tripping on acid ...

OpenStudy (anonymous):

Pythagorean Theorem is the lesson this derives from? I don't even know The Law of Cosines..?

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

sin = opposite over hypotenuse tan = opposite over adjacent cos = adjacent over hypotenuse

OpenStudy (anonymous):

you would use cosine =)

OpenStudy (anonymous):

I don't use that in my Geometry class? In fact I've never even been introduced to that.

OpenStudy (anonymous):

5/8.24621125124

OpenStudy (anonymous):

you have no other angles to go off of =)

OpenStudy (anonymous):

so therefor you have no other possible way of knowing =)

OpenStudy (anonymous):

Except for the Pythagorean Theorem. \[a^2+b^2=c^2\]

OpenStudy (anonymous):

yeah you wouldnt use that here =)

OpenStudy (anonymous):

If a^2+b^2=c^2, its a right triangle.

OpenStudy (anonymous):

you have no angles =)

OpenStudy (anonymous):

If they are less than it's acute. Greater than it's obtuse.

OpenStudy (anonymous):

and you wouldnt use cosine btw

OpenStudy (anonymous):

wait yeh you would lol

OpenStudy (anonymous):

3^2+5^2 does not equal 2 radical 17

OpenStudy (anonymous):

So what does C have to be to make them equal?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

you answered your own question for the 1st part of thiss question =)

OpenStudy (anonymous):

9+25= 34

OpenStudy (anonymous):

"What would C have to be to make it a right triangle?" Was my question.

OpenStudy (anonymous):

So it would be radical 34?

OpenStudy (anonymous):

yes =)

OpenStudy (anonymous):

Thanks. c:

OpenStudy (anonymous):

if a^2 +b^2 = C^2

OpenStudy (anonymous):

then yeah you would need the radical other than that it would be 5.83095189485^2

OpenStudy (anonymous):

you will eventually learn soh cah toa =)

OpenStudy (anonymous):

you can find angle measurements with that

OpenStudy (anonymous):

thats what i thought you were doing lol =)